The sum of three consecutive integers is 33 more than the least of the integers. Find the integers.
Found 2 solutions by elima, bucky:
Answer by elima(1433)

The sum of three consecutive integers is 33 more than the least of the integers. Find the integers.
x=first integer
x+1=second integer
x+2= third integer
========================
x+x+1+x+2=x+2+33
3x+3=x+35
3x-x=35-3
2x=32
x=16
so the integers are;
16,17,18
:)

Answer by bucky(2189)

The three consecutive integers can be represented as:
.
x
x+1 and
x+2
.
The sum of these three is 3x + 3
.
If you take 33 away from this sum, it is to equal the least of the integers which is x.
So
.
3x + 3 - 33 = x
.
Subtract x from both sides and you get:
.
2x + 3 - 33 = 0
.
Now combine the +3 and -33 and you have:
.
2x - 30 = 0
.
Add 30 to both sides:
.
2x = 30
.
Solve for x by dividing both sides by 2 to get:
.
x = 15
.
So the answer is that the three integers are 15, 16, and 17.
.